two (2) nickel coins with identical charges are separated by 1.5 m. If the forces exerted by the coins to one another are both 2N, what is the charge on ech coin?


1
Expert's answer
2022-03-22T15:47:23-0400

We know that

"F=\frac{kq_1q_2}{r^2}"

"F=\frac{kq^2}{r^2}"

"q=\sqrt{\frac{Fr^2}{k}}"

"q_1=q_2=q=\sqrt{\frac{2\times 1.5^2}{9\times10^9}}=22.36\times10^{-6}C"